# Benchmarking Content-Based Puzzle Solvers on Corrupted Jigsaw Puzzles

### v1

@@ Abstract
Keywords:
Jigsaw Puzzle Solvers Robustness Image Processing
{NoHyper}Accepted at ICIAP 2025.
This preprint has not undergone peer review or any post-submission improvements or corrections.
@@ 1 Introduction
Real-world combinatorial optimization problems are NP-hard, complex, and time-consuming to solve. In addition to these challenges, automatic solvers must handle noisy data and non-optimal surrogate objectives, making robustness a critical requirement. One category of these real-world problems includes the reassembly of objects such as shredded documents [7] or fragmented historical artworks [9, 10]. This challenge is in close relationship to solving jigsaw puzzles as both are defined similarly: Given a set of pieces, the goal is to reconstruct the original image or object, from which the pieces were created [6]. Throughout the years, many automated solvers for jigsaw puzzles were developed utilizing both heuristic and deep learning methods [6] with the hope of also being applicable to real-world problems.
@@
Despite significant advancements, a recent study [17] demonstrates that nearly all existing jigsaw puzzle solvers are unsuitable to work on realistic archaeological fragments and those that can be applied perform poorly. This is because the fragments are eroded, irregularly shaped due to fractures, vary in size, or are missing due to weathering or human-induced damage. Such characteristics of a real-world archaeological puzzle are often not present in the datasets that are used to develop jigsaw puzzle solvers. Therefore, the authors identify a research gap between existing computational methods and the automated reconstruction of real-world artefacts.
@@
To bridge this gap, it is essential to evaluate the robustness of existing solvers by systematically assessing their ability to handle these real-world challenges. Since isolating these characteristics from an archaeological dataset is impossible, a practical approach is to introduce controlled corruptions to standard jigsaw puzzles mimicking the complexities of real-world artefacts. These types of corruptions can be categorized into two groups: The first group involves modifications to the shape of a puzzle piece, such as irregular contours, multiple edges, or varying piece sizes [3, 4]. The existing approaches to these challenges focus on fitting together the contours of neighbouring pieces, while mostly disregarding the pictorial content of the fragments. The second group includes corruptions that alter the visual content of a puzzle, such as erosion of texture and colour, the degradation of the edges of a piece, or the complete absence of multiple pieces. Most content-based jigsaw puzzle solvers are designed to reassemble uncorrupted, square-sized pieces [6] and their robustness to such modifications remains largely unexplored.
@@
In this paper, we focus on content-modifying corruptions and systematically benchmark the State-Of-The-Art (SOTA) content-based jigsaw puzzle solvers to investigate their robustness to different types of corruptions. Using the PuzzleWikiArts dataset [14] we introduce three types of corruptions to simulate real-world fragment characteristics. We then evaluate the performance of selected solvers, comparing heuristic methods and deep learning models, to determine their effectiveness and limitations. Furthermore, we explore whether learning-based methods can be adapted to handle corruption more effectively. Based on our structured investigation, we derive suggestions for the most promising research directions for developing automated methods capable of reconstructing historical artefacts.
@@ 2 Related Work
When reviewing existing research, we focus on 2D content-based image puzzles with square-sized, non-rotated pieces, also known as Type 1 jigsaw puzzles [1], as it is the most frequently studied type in the literature [6]. We identify the SOTA puzzle solvers to test them in our benchmark by examining the comparative investigation by Giuliari et al. [2], which compare three heuristic solvers and four deep learning models, and Son et al. [13], which compiled the results reported for nine heuristic solvers.
@@ 2.1 Content-Based Heuristic Jigsaw Puzzle Solvers
Over the years, many heuristic algorithms have been proposed to automatically solve Type 1 jigsaw puzzles. Most of these methods consist of two key components: a metric to compute the pairwise compatibility between neighbouring puzzle pieces and a strategy to assemble the puzzle based on this metric. One of those pairwise compatibility measures, introduced by Gallagher [1], is the Mahalanobis Gradient Compatibility (MGC). It evaluates the local gradients at the boundaries of the jigsaw puzzle pieces, based on the assumption that neighbouring pieces should exhibit similar gradient distributions at their edges. Gallagher shows that MGC significantly outperforms previous compatibility metrics which rely on pixel intensity differences at piece boundaries. This is also confirmed by Giuliari et al. [2] and Son et al. [13]. Gallagher combines the MGC with a greedy, tree-based reassembly algorithm to solve different types of jigsaw puzzles.
The pairwise compatibility metric by Paikin and Tal [8] assesses the $L1$ norm of the boundary pixels between all pieces as a measurement of dissimilarity. Their deterministic heuristic identifies pairs of pieces with the smallest and most unique differences, as these are likely to be neighbours, and uses them as starting points to greedily reconstruct the original image. Their method achieves high reconstruction accuracies, outperforming Gallagher’s solver [13, 2].
Yu et al. [12] utilize MGC as their pairwise compatibility metric and combine it with a linear program, which exploits all pairwise matches simultaneously. According to the comparison by Son et al. [13], Yu et al.’s approach achieves the highest performance on two out of five jigsaw puzzle datasets and the second-highest performance on the remaining three datasets.
@@ 2.2 Content-Based Deep Learning Jigsaw Puzzle Solvers
In recent years, many researchers tried to apply deep learning models to the jigsaw puzzle problem. These methods follow a two-step process: first, visual features are extracted from the individual image pieces by a neural network. In the second step, the original permutation or the absolute positions of the puzzle pieces are predicted based on the processed features. Therefore, deep learning methods have the advantage that they can learn the reassembly task by leveraging both structure- and colour-based features across the entire image [6]. However, earlier deep learning approaches were limited to small puzzle sizes or a fixed number of possible input permutations.
Talon et al. [14, 15] are the first who tested their model on puzzles up to $12\times 12$ pieces but their performance is still behind those of the best heuristics. As one of the few papers in this field, they test how different approaches handle missing pieces, Gaussian noise, or missing borders in the puzzle. They report a performance drop for all investigated solvers but do not analyse the results further.
A recent paper by Giuliari et al. [2] introduced Positional Diffusion, a method based on Diffusion Probabilistic Models that is designed to solve ordering problems. For image puzzles, the method learns to predict the positions of the pieces in a normalized coordinate system by perturbing their position vectors during the diffusion process. The authors report that their method outperforms other solvers, for example, Paikin and Tal’s approach, across most of the tested puzzle sizes. Notably, one of the compared deep learning solvers that is also developed by the same authors is a Transformer model that likewise predicts the pieces’ coordinates but without the diffusion process.
In summary, research on applying deep learning models to the jigsaw puzzle problem is still limited, and only a few approaches achieve performance comparable to existing heuristic solvers. As a result, we include two deep learning models in our benchmark: the Transformer-based model and Positional Diffusion, both by Giuliari et al.  [^1] Implementations: Gallagher, Yu et al. missing, Paikin and Tal, Positional Diffusion
@@ 3 Methodology
Figure 1: Example of a $12\times 12$ puzzle modified by each of the three corruption types. The original image is taken from the PuzzleWikiArts dataset [14].
This chapter presents our approach to benchmark the five selected, SOTA content-based jigsaw puzzle solvers on corrupted puzzles to determine their robustness. To test the performance of the solvers, we are analysing two established metrics [1]. The Direct Comparison Metric measures the percentage of pieces correctly placed in their absolute positions within the puzzle. The Perfect Reconstruction Metric is the percentage of jigsaw puzzles from a dataset that could be correctly reassembled into the original image. For our experiments, we define three types of corruptions inspired by real-world archaeological puzzles and describe their systematic application to unmodified jigsaw puzzles: missing pieces, eroded edges, and eroded contents.
@@ 3.1.1 Missing Pieces
The first type of corruption includes missing pieces. Therefore, a percentage of the total number of pieces is removed from the puzzle and the solvers are tasked with reconstructing the original image as accurately as possible using the remaining pieces. The missing pieces are randomly selected, which can lead to holes in the image, as shown in Fig. 1(b) if the proportion is low. At higher percentages, solvers may need to assemble disconnected components of the image, further increasing the complexity of the task. In each experiment, the percentage of missing pieces is gradually increased and we limit the experiments to a maximum of 50%. Since all three heuristic solvers do not inherently support missing pieces, the selected pieces are replaced with black image patches of the same shape, with pixel values set to zero. In comparison, it is valid to remove the selected missing pieces from the input for both investigated deep learning models as they predict the absolute position of each puzzle piece. To assess the robustness of the puzzle solvers against this type of corruption, we analyse the two performance metrics. Since only the correct placement of the non-missing pieces should be considered, the computation of these metrics is adapted accordingly.
@@ 3.1.2 Eroded Edges
To simulate eroded edges on square-piece jigsaw puzzles in our benchmark, we adopt the approach of Huroyan et al. [5], who use this type of corruption to compare three heuristic solvers. Specifically, we randomly select any of the four edges of a puzzle piece with a fixed probability. For each selected edge, the two outermost pixel rows or columns are replaced with new values, which are uniformly sampled from the entire image. An example of this corruption is shown in Fig. 1(c). While Huroyan et al. examine eroded edges with probabilities up to 14%, we extend this range by gradually increasing the probability up to 50%. At this maximum level, an average of two edges per piece are affected, severely limiting the puzzle solvers’ ability to compare neighbouring edges and reconstruct the original image.
@@ 3.1.3 Eroded Contents
The third category of corruption includes erosion suffered to the pictorial contents. As the causes and the severity of damage to an artwork differ for each real-world example, it is difficult to simulate all possibilities of decay. Therefore, we focus on four general effects that can impact historical paintings which are exposed to erosion, specifically loss in saturation, contrast and brightness as well as the complete loss of paint through flaking [11].
To control the amount of erosion to the contents of a jigsaw puzzle, we introduce the erosion factor $\mathcal{E}\in[0,100]$ (in %) which controls the corruption of the pieces in two ways. First, we assume that in a realistic fragmented artwork, each fragment is affected differently by erosion, meaning that some parts could be unharmed while others are severely damaged. $\mathcal{E}$ controls the proportion of pieces that are modified by each of the four erosion effects. This means that a piece in the jigsaw puzzle can be either unmodified or modified by one or a combination of multiple corruption types. Secondly, $\mathcal{E}$ also controls the severity of the individual erosion effect. If a piece is selected for corruption, the proposition of the effect is sampled from a discrete list using a normal distribution centred around $\mathcal{E}$. As a result, the severity of the erosion is generally increased with higher $\mathcal{E}$, while each piece of a puzzle is affected differently by the four effects. A visual example of this type of corruption can be seen in Fig. 1(d).
@@ 3.2 Content-Based Jigsaw Puzzle Dataset
We apply the selected corruptions on the PuzzleWikiArts dataset [14]. It is derived from WikiArts [16] and includes 63 thousand images of paintings, featuring a wide variety of contents and artistic styles. The images are commonly split into puzzles of $6\times 6$ up to $12\times 12$ pieces with each piece consisting of $32\times 32$ pixels. For testing, the four puzzle sizes are uniformly distributed across the test set, resulting in approximately 3125 images per puzzle size.
@@ 3.3 Experimental Design
For all solvers, we use their standard implementations, with adjustments made only to the data loading and evaluation procedures to accommodate the specific types of corruption. Before benchmarking the solvers on corrupted puzzles, the SOTA solvers are first tested on standard Type 1 jigsaw puzzles without any corruption. This step is done to reproduce the performance of the four selected approaches as reported in [2]. Additionally, we compare the solvers based on their overall performance across the two evaluation metrics. This serves as a baseline before proceeding with the subsequent experiments, where we investigate the robustness of the jigsaw puzzle solvers towards missing pieces (Experiment 1), eroded edges (Experiment 2) and eroded contents (Experiment 3).
@@ 4.1 Baseline: Standard Jigsaw Puzzles
To establish a baseline for the subsequent experiments, we evaluate all five selected jigsaw puzzle solvers on standard Type 1 puzzles. The results are presented in Tab. 1. Comparing our results with those of Giuliari et al. [2] on the Direct Comparison Metric, we successfully reproduce the performance of three out of four solvers. Only for the solver by Paikin and Tal, we observe a significant divergence, with an average performance gap of 13.85% compared to the previously reported results. The reason for this discrepancy remains unclear, especially since the exact implementation used in the previous study is not specified.
@@ 4.2 Experiment 1: Corrupting Jigsaw Puzzles with Missing Pieces
In general, the performance of all jigsaw puzzle solvers declines as the number of missing pieces increases. This trend is evident in the evaluation metrics across all four puzzle sizes, as shown in Fig. 2. Among the heuristics, Yu et al.’s solver generally achieves the highest performance, followed by the methods of Paikin and Tal and then Gallagher. The Transformer and Positional Diffusion base models also perform worse than Yu et al.’s heuristic.
Fine-tuning the deep learning models on puzzles with missing pieces significantly improves their performance. Although the Transformer model still lags behind, Positional Diffusion outperforms Yu et al.’s heuristic on smaller puzzles with up to $8\times 8$ pieces in the Direct Comparison Metric. However, for larger $12\times 12$ puzzles, Yu et al.’s solver remains superior when 30% or fewer pieces are missing. Additionally, the heuristic reconstructs more jigsaw puzzles perfectly than Positional Diffusion for puzzles of size $8\times 8$ or larger. The visual example in Tab. 2 presents the prediction of the five puzzle solvers on a jigsaw puzzle with 10% missing pieces. In this example, the three heuristics and Positional Diffusion can reconstruct the original image perfectly, while the fine-tuned Transformer model misplaces multiple pieces.

### v2

@@ Abstract
{NoHyper}This preprint has not undergone peer review or any post-submission improvements or corrections. The Version of Record of this contribution is published in
ICIAP 2025, Lecture Notes in Computer Science 16167 (2026) pp 286-298,
and is available online at https://doi.org/10.1007/978-3-032-10185-3_23
@@ 1 Introduction
Real-world combinatorial optimization problems are NP-hard, complex, and time-consuming to solve. In addition to these challenges, automatic solvers must handle noisy data and non-optimal surrogate objectives, making robustness a critical requirement. One category of these real-world problems includes the reassembly of objects such as shredded documents [marques_reconstructing_2009] or fragmented historical artworks [papaodysseus_contour-shape_2002, pintus_survey_2016]. This challenge is in close relationship to solving jigsaw puzzles as both are defined similarly: Given a set of pieces, the goal is to reconstruct the original image or object, from which the pieces were created [markaki_jigsaw_2023]. Throughout the years, many automated solvers for jigsaw puzzles were developed utilizing both heuristic and deep learning methods [markaki_jigsaw_2023] with the hope of also being applicable to real-world problems.
@@
Despite significant advancements, a recent study [tsesmelis_re-assembling_2024] demonstrates that nearly all existing jigsaw puzzle solvers are unsuitable to work on realistic archaeological fragments and those that can be applied perform poorly. This is because the fragments are eroded, irregularly shaped due to fractures, vary in size, or are missing due to weathering or human-induced damage. Such characteristics of a real-world archaeological puzzle are often not present in the datasets that are used to develop jigsaw puzzle solvers. Therefore, the authors identify a research gap between existing computational methods and the automated reconstruction of real-world artefacts.
@@
To bridge this gap, it is essential to evaluate the robustness of existing solvers by systematically assessing their ability to handle these real-world challenges. Since isolating these characteristics from an archaeological dataset is impossible, a practical approach is to introduce controlled corruptions to standard jigsaw puzzles mimicking the complexities of real-world artefacts. These types of corruptions can be categorized into two groups: The first group involves modifications to the shape of a puzzle piece, such as irregular contours, multiple edges, or varying piece sizes [harel_crossing_2021, harel_pictorial_2024]. The existing approaches to these challenges focus on fitting together the contours of neighbouring pieces, while mostly disregarding the pictorial content of the fragments. The second group includes corruptions that alter the visual content of a puzzle, such as erosion of texture and colour, the degradation of the edges of a piece, or the complete absence of multiple pieces. Most content-based jigsaw puzzle solvers are designed to reassemble uncorrupted, square-sized pieces [markaki_jigsaw_2023] and their robustness to such modifications remains largely unexplored.
@@
In this paper, we focus on content-modifying corruptions and systematically benchmark the State-Of-The-Art (SOTA) content-based jigsaw puzzle solvers to investigate their robustness to different types of corruptions. Using the PuzzleWikiArts dataset [talon_ganzzle_2022] we introduce three types of corruptions to simulate real-world fragment characteristics. We then evaluate the performance of selected solvers, comparing heuristic methods and deep learning models, to determine their effectiveness and limitations. Furthermore, we explore whether learning-based methods can be adapted to handle corruption more effectively. Based on our structured investigation, we derive suggestions for the most promising research directions for developing automated methods capable of reconstructing historical artefacts.
@@ 2 Related Work
When reviewing existing research, we focus on 2D content-based image puzzles with square-sized, non-rotated pieces, also known as Type 1 jigsaw puzzles [gallagher_jigsaw_2012], as it is the most frequently studied type in the literature [markaki_jigsaw_2023]. We identify the SOTA puzzle solvers to test them in our benchmark by examining the comparative investigation by Giuliari et al. [giuliari_positional_2024], which compare three heuristic solvers and four deep learning models, and Son et al. [son_solving_2019], which compiled the results reported for nine heuristic solvers.
@@ 2.1 Content-Based Heuristic Jigsaw Puzzle Solvers
Over the years, many heuristic algorithms have been proposed to automatically solve Type 1 jigsaw puzzles. Most of these methods consist of two key components: a metric to compute the pairwise compatibility between neighbouring puzzle pieces and a strategy to assemble the puzzle based on this metric. One of those pairwise compatibility measures, introduced by Gallagher [gallagher_jigsaw_2012], is the Mahalanobis Gradient Compatibility (MGC). It evaluates the local gradients at the boundaries of the jigsaw puzzle pieces, based on the assumption that neighbouring pieces should exhibit similar gradient distributions at their edges. Gallagher shows that MGC significantly outperforms previous compatibility metrics which rely on pixel intensity differences at piece boundaries. This is also confirmed by Giuliari et al. [giuliari_positional_2024] and Son et al. [son_solving_2019]. Gallagher combines the MGC with a greedy, tree-based reassembly algorithm to solve different types of jigsaw puzzles.
The pairwise compatibility metric by Paikin and Tal [paikin_solving_2015] assesses the $L1$ norm of the boundary pixels between all pieces as a measurement of dissimilarity. Their deterministic heuristic identifies pairs of pieces with the smallest and most unique differences, as these are likely to be neighbours, and uses them as starting points to greedily reconstruct the original image. Their method achieves high reconstruction accuracies, outperforming Gallagher’s solver [son_solving_2019, giuliari_positional_2024].
Yu et al. [yu_solving_2015] utilize MGC as their pairwise compatibility metric and combine it with a linear program, which exploits all pairwise matches simultaneously. According to the comparison by Son et al. [son_solving_2019], Yu et al.’s approach achieves the highest performance on two out of five jigsaw puzzle datasets and the second-highest performance on the remaining three datasets.
@@ 2.2 Content-Based Deep Learning Jigsaw Puzzle Solvers
In recent years, many researchers tried to apply deep learning models to the jigsaw puzzle problem. These methods follow a two-step process: first, visual features are extracted from the individual image pieces by a neural network. In the second step, the original permutation or the absolute positions of the puzzle pieces are predicted based on the processed features. Therefore, deep learning methods have the advantage that they can learn the reassembly task by leveraging both structure- and colour-based features across the entire image [markaki_jigsaw_2023]. However, earlier deep learning approaches were limited to small puzzle sizes or a fixed number of possible input permutations.
Talon et al. [talon_ganzzle_2022, talon_ganzzle_2025] are the first who tested their model on puzzles up to $12\times 12$ pieces but their performance is still behind those of the best heuristics. As one of the few papers in this field, they test how different approaches handle missing pieces, Gaussian noise, or missing borders in the puzzle. They report a performance drop for all investigated solvers but do not analyse the results further.
A recent paper by Giuliari et al. [giuliari_positional_2024] introduced Positional Diffusion, a method based on Diffusion Probabilistic Models that is designed to solve ordering problems. For image puzzles, the method learns to predict the positions of the pieces in a normalized coordinate system by perturbing their position vectors during the diffusion process. The authors report that their method outperforms other solvers, for example, Paikin and Tal’s approach, across most of the tested puzzle sizes. Notably, one of the compared deep learning solvers that is also developed by the same authors is a Transformer model that likewise predicts the pieces’ coordinates but without the diffusion process.
In summary, research on applying deep learning models to the jigsaw puzzle problem is still limited, and only a few approaches achieve performance comparable to existing heuristic solvers. As a result, we include two deep learning models in our benchmark: the Transformer-based model and Positional Diffusion, both by Giuliari et al.  [^1] Implementations: Gallagher, Yu et al., Paikin and Tal, Positional Diffusion
@@ 3 Methodology
Figure 1: Example of a $12\times 12$ puzzle modified by each of the three corruption types. The original image is taken from the PuzzleWikiArts dataset [talon_ganzzle_2022].
This chapter presents our approach to benchmark the five selected, SOTA content-based jigsaw puzzle solvers on corrupted puzzles to determine their robustness. To test the performance of the solvers, we are analysing two established metrics [gallagher_jigsaw_2012]. The Direct Comparison Metric measures the percentage of pieces correctly placed in their absolute positions within the puzzle. The Perfect Reconstruction Metric is the percentage of jigsaw puzzles from a dataset that could be correctly reassembled into the original image. For our experiments, we define three types of corruptions inspired by real-world archaeological puzzles and describe their systematic application to unmodified jigsaw puzzles: missing pieces, eroded edges, and eroded contents.
@@ 3.1.1 Missing Pieces
The first type of corruption includes missing pieces. Therefore, a percentage of the total number of pieces is removed from the puzzle and the solvers are tasked with reconstructing the original image as accurately as possible using the remaining pieces. The missing pieces are randomly selected, which can lead to holes in the image, as shown in Fig.˜1(b) if the proportion is low. At higher percentages, solvers may need to assemble disconnected components of the image, further increasing the complexity of the task. In each experiment, the percentage of missing pieces is gradually increased and we limit the experiments to a maximum of 50%. Since all three heuristic solvers do not inherently support missing pieces, the selected pieces are replaced with black image patches of the same shape, with pixel values set to zero. In comparison, it is valid to remove the selected missing pieces from the input for both investigated deep learning models as they predict the absolute position of each puzzle piece. To assess the robustness of the puzzle solvers against this type of corruption, we analyse the two performance metrics. Since only the correct placement of the non-missing pieces should be considered, the computation of these metrics is adapted accordingly.
@@ 3.1.2 Eroded Edges
To simulate eroded edges on square-piece jigsaw puzzles in our benchmark, we adopt the approach of Huroyan et al. [huroyan_solving_2020], who use this type of corruption to compare three heuristic solvers. Specifically, we randomly select any of the four edges of a puzzle piece with a fixed probability. For each selected edge, the two outermost pixel rows or columns are replaced with new values, which are uniformly sampled from the entire image. An example of this corruption is shown in Fig.˜1(c). While Huroyan et al. examine eroded edges with probabilities up to 14%, we extend this range by gradually increasing the probability up to 50%. At this maximum level, an average of two edges per piece are affected, severely limiting the puzzle solvers’ ability to compare neighbouring edges and reconstruct the original image.
@@ 3.1.3 Eroded Contents
The third category of corruption includes erosion suffered to the pictorial contents. As the causes and the severity of damage to an artwork differ for each real-world example, it is difficult to simulate all possibilities of decay. Therefore, we focus on four general effects that can impact historical paintings which are exposed to erosion, specifically loss in saturation, contrast and brightness as well as the complete loss of paint through flaking [podany_art_2023].
To control the amount of erosion to the contents of a jigsaw puzzle, we introduce the erosion factor $\mathcal{E}\in[0,100]$ (in %) which controls the corruption of the pieces in two ways. First, we assume that in a realistic fragmented artwork, each fragment is affected differently by erosion, meaning that some parts could be unharmed while others are severely damaged. $\mathcal{E}$ controls the proportion of pieces that are modified by each of the four erosion effects. This means that a piece in the jigsaw puzzle can be either unmodified or modified by one or a combination of multiple corruption types. Secondly, $\mathcal{E}$ also controls the severity of the individual erosion effect. If a piece is selected for corruption, the proposition of the effect is sampled from a discrete list using a normal distribution centred around $\mathcal{E}$. As a result, the severity of the erosion is generally increased with higher $\mathcal{E}$, while each piece of a puzzle is affected differently by the four effects. A visual example of this type of corruption can be seen in Fig.˜1(d).
@@ 3.2 Content-Based Jigsaw Puzzle Dataset
We apply the selected corruptions on the PuzzleWikiArts dataset [talon_ganzzle_2022]. It is derived from WikiArts [tan_improved_2019] and includes 63 thousand images of paintings, featuring a wide variety of contents and artistic styles. The images are commonly split into puzzles of $6\times 6$ up to $12\times 12$ pieces with each piece consisting of $32\times 32$ pixels. For testing, the four puzzle sizes are uniformly distributed across the test set, resulting in approximately 3125 images per puzzle size.
@@ 3.3 Experimental Design
For all solvers, we use their standard implementations, with adjustments made only to the data loading and evaluation procedures to accommodate the specific types of corruption. Before benchmarking the solvers on corrupted puzzles, the SOTA solvers are first tested on standard Type 1 jigsaw puzzles without any corruption. This step is done to reproduce the performance of the four selected approaches as reported in [giuliari_positional_2024]. Additionally, we compare the solvers based on their overall performance across the two evaluation metrics. This serves as a baseline before proceeding with the subsequent experiments, where we investigate the robustness of the jigsaw puzzle solvers towards missing pieces (Experiment 1), eroded edges (Experiment 2) and eroded contents (Experiment 3).
@@ 4.1 Baseline: Standard Jigsaw Puzzles
To establish a baseline for the subsequent experiments, we evaluate all five selected jigsaw puzzle solvers on standard Type 1 puzzles. The results are presented in Tab.˜1. Comparing our results with those of Giuliari et al. [giuliari_positional_2024] on the Direct Comparison Metric, we successfully reproduce the performance of three out of four solvers. Only for the solver by Paikin and Tal, we observe a significant divergence, with an average performance gap of 13.85% compared to the previously reported results. The reason for this discrepancy remains unclear, especially since the exact implementation used in the previous study is not specified.
@@ 4.2 Experiment 1: Corrupting Jigsaw Puzzles with Missing Pieces
In general, the performance of all jigsaw puzzle solvers declines as the number of missing pieces increases. This trend is evident in the evaluation metrics across all four puzzle sizes, as shown in Fig.˜2. Among the heuristics, Yu et al.’s solver generally achieves the highest performance, followed by the methods of Paikin and Tal and then Gallagher. The Transformer and Positional Diffusion base models also perform worse than Yu et al.’s heuristic.
Fine-tuning the deep learning models on puzzles with missing pieces significantly improves their performance. Although the Transformer model still lags behind, Positional Diffusion outperforms Yu et al.’s heuristic on smaller puzzles with up to $8\times 8$ pieces in the Direct Comparison Metric. However, for larger $12\times 12$ puzzles, Yu et al.’s solver remains superior when 30% or fewer pieces are missing. Additionally, the heuristic reconstructs more jigsaw puzzles perfectly than Positional Diffusion for puzzles of size $8\times 8$ or larger. The visual example in Tab.˜2 presents the prediction of the five puzzle solvers on a jigsaw puzzle with 10% missing pieces. In this example, the three heuristics and Positional Diffusion can reconstruct the original image perfectly, while the fine-tuned Transformer model misplaces multiple pieces.